# Efficient Input-Constrained Impulsive Optimal Control of Linear Systems with Application to Spacecraft Relative Motion

### v1

@@ Abstract
This work presents a novel algorithm for impulsive optimal control of linear time-varying systems with the inclusion of input magnitude constraints. Impulsive optimal control problems, where the optimal input solution is a sum of delta functions, are typically formulated as an optimization over a normed function space subject to integral equality constraints and can be efficiently solved for linear time-varying systems in their dual formulation. In this dual setting, the problem takes the form of a semi-infinite program which is readily solvable in online scenarios for constructing maneuver plans. This work augments the approach with the inclusion of magnitude constraints on the input over time windows of interest, which is shown to preserve the impulsive nature of the optimal solution and enable efficient solution procedures via semi-infinite programming. The resulting algorithm is demonstrated on the highly relevant problem of relative motion control of spacecraft in Low Earth Orbit (LEO) and compared to several other proposed solutions from the literature.
@@ I Introduction
The problem of minimizing the norm of a control input function space of a linear time-varying system with boundary conditions has received significant attention due to its practical application to many systems of interest. Such optimal control problems arise in many fields, including astrodynamics [1], epidemiology [2], robotics [3], and economics [4]. Despite its widespread applicability, impulsive optimal control remains challenging for many applications of interest, which may involve complex state and input constraints and require efficient and reliable solution procedures. For example, in the context of spacecraft relative motion control, for which computational resources are limited, safety is essential, and algorithmic reliability is crucial, few existing approaches are adequate.
Approaches to solving impulsive optimal control problems are numerous, and include direct methods, such as sequential convex programming [5], indirect methods, leveraging moment problems [6], reachable set theory [7], or primer-vector theory [1], and other techniques, like particle swarm optimization [8] and analytical methods [9]. In direct methods, the dynamics and cost are discretized over a time grid of interest and solved as finite-dimensional optimization problem using tools like convex optimization [10]. These techniques have a tendency to introduce a large number of optimization variables and constraints and can produce suboptimal policies in the case of coarse gridding. Remedies to this issue include having impulse times as optimization variables [5], but this results in a non-linear program that must be solved through techniques like sequential convex programming [11], which may not necessarily provide convergence guarantees [12]. Despite this, these techniques are highly desirable due to their generality and ability to handle input and state constraints. In contrast to direct methods, analytical methods directly solve optimality conditions to derive analytical expressions for control signals but suffer from limited applicability as the solutions can be complicated and specific to a given system. However, they are highly reliable and efficient, and have seen success in the context of spacecraft control [9]. Finally, other techniques like particle swarm optimization [8] and genetic algorithms [13] have also been studied but are typically not well equipped for efficiency and generalizability.
Indirect methods, in comparison, involve formulating optimality conditions on the continuous-time problem, and include moment-based formulations, reachable set theory, and primer-vector theory approaches, which provide different perspectives to the problem and yet all arrive at equivalent optimality conditions. Primer-vector theory, originally coined by Lawden [14] in 1963, derives first order necessary conditions of optimality which can be derived from Pontryagin’s Maximum Principle. The moment problem dates back to the 1960s, when Neustadt [6] showed that the optimal control problem over normed integrable function spaces with integral equality constraints can be embedded in a conjugate space and that the corresponding function space is a sum of delta functions. In the 1970s, Gilbert [7] approached the problem from the reachable set theory perspective, formulating necessary conditions for optimality in terms of spacecraft reachable sets to derive efficient algorithms for fuel-optimal spacecraft rendezvous.
@@
Early algorithmic frameworks for leveraging the optimality conditions derived from these indirect methods are proposed in Gilbert [7], which employed an early numerical algorithm based on quadratic programming [15] to efficiently solve impulsive rendezvous problems. Other approaches involve simplified techniques to solve the necessary conditions to suboptimality [16] or applying mixed iterative schemes to solving the polynomial necessary and sufficient conditions of optimality [17]. In 2013, Arzelier [18] connected the moment problem theorized by Neustadt [6] to a semi-infinite convex optimization scheme [19] to produce an efficient, convergent, and reliable strategy to solving the linearized impulsive optimal control problem. The discretization strategy proposed in this work has served as the foundation for many impulsive optimal control studies for spacecraft relative motion control, such as [20], which connected the technique to reachable set theory and extended the technique to time varying cost functions, and [21], which proposed an extended algorithm to handle state constraints and input constraints. This semi-infinite program (SIP) approach has seen great success in the context of spacecraft rendezvous and relative motion control due to its efficiency and reliability, with heritage on CubeSat missions like VISORS [22].
@@
Despite the unilateral utility of these theoretical and algorithmic frameworks, the applicability of indirect methods to constrained problems remains limited. With respect to spacecraft rendezvous, a major constraint of interest involves bounding the maximum impulse that a spacecraft can execute at a given time. All satellites, such as those of the VISORS mission, have propulsion systems that naturally constrain the amount of $\Delta V$ that can be applied to the system at a given time, necessitating the use of maneuver splitting algorithms or other heuristic approaches [22] which lead to suboptimal maneuver plans and generally increased computational burden. More broadly, it is often practical to distribute $\Delta V$ expenditure for spacecraft relative motion irrespective of propulsion constraints to reduce the aggressiveness of maneuver plans. In the case of rendezvous over 10s of kilometers, naively applying a scheme without the consideration of impulse magnitude can result in maneuver plans that are unsafe, unrealizable, or otherwise undesirable.
@@
To address this theoretical and algorithmic gap, this paper introduces impulse magnitude constraints by applying inequality constraints over the integral of input normed function spaces. It is shown that the resulting formulation preserves the impulsive quality of the solution, and can be easily extended to semi-infinite programming techniques to produce an efficient and reliable solution procedure. The algorithm is demonstrated for several optimal control problems of interest in spacecraft relative motion control and compared with existing approaches.
@@ II Problem Background
where $\bm{x}_{0}\in\mathbb{R}^{n_{x}}$ and $\bm{x}_{f}\in\mathbb{R}^{n_{x}}$ are fixed initial and terminal constraints. The set of input functionals that satisfy the system dynamics and boundary conditions can be expressed as an equality constraint in terms of the state transition matrix via
@@
Problem 1 is a moment problem of the same form given in Neustadt [6] whose optimal solution is a sum of delta functions, i.e. the input is an atomic measure. In its primal form, this problem is difficult to solve due to its infinite dimensional nature. The dual form of the problem can be derived from the Lagrangian, which takes the form
$\mathcal{L}(\bm{u}(\cdot),$ $\bm{\lambda}_{f})=\bm{\lambda_{f}}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})$ (5)
$+\int_{t_{0}}^{t_{f}}f(\bm{u}(\tau))-\bm{\lambda}_{f}^{\intercal}\Phi(t_{f},\tau)B(\tau)u(\tau)\text{d}\tau,$
where $\bm{\lambda}_{f}$ is a Lagrangian multiplier associated with the equality constraint in Equation 3 and has a dual function given by
$g(\bm{\lambda}_{f})=\inf_{\bm{u}(\cdot)}\mathcal{L}(\bm{u}(\cdot),\bm{\lambda}_{f})$ (6)
[…]
Here, $f^{\circ}(\cdot):\mathbb{R}^{n_{u}}\to\mathbb{R}^{+}$ is the dual norm of $f$, i.e. $f^{\circ}(\bm{y})=\sup_{f(\bm{x})\leq 1}\bm{x}^{\intercal}\bm{y}$. The dual form of the problem is arrived to by the program formed by maximizing the dual function, i.e.
@@
$\mathcal{D}=\max_{\bm{\lambda}_{f}}$ $\bm{\lambda}_{f}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})$ (7)
@@
Problem 7 consists of maximizing an affine function subject to a semi-infinite constraint over time over the norm of the Lagrangian multiplier. As such, it is convex.
In contrast to the primal problem, which consists of minimization of a function space, and is therefore infinite dimensional, the dual problem is an optimization problem over a finite number of variables subject to an infinite constraint space. The dual form of this problem is in fact easier to solve in this form, as Arzelier [18] showed, through leveraging standard discretization methods for semi-infinite programs.
The moment problem is connected to both primer-vector theory and reachable set theory approaches through its dual form. Indeed, $\bm{\lambda}_{v}(t)\triangleq B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}$ is in fact the classic primer-vector [14] which satisfies $\dot{\bm{\lambda}}(t)=-A^{\intercal}(t)\bm{\lambda}(t),\bm{\lambda}_{v}(t)=B^{\intercal}(t)\bm{\lambda}(t)$. In this sense, the Lagrangian multiplier $\bm{\lambda}_{f}$ introduced in Equation 5 is simply the costate at the terminal time and the semi-infinite constraint in Problem 7 is equivalent to that derived in Prussing [1] ( $\lambda_{v}^{\intercal}\lambda_{v}\leq 1$ in the case of the $L_{2}$ norm). The same comparison can be made for the more geometric interpretation of reachable set theory [7], which is formulated in terms of support functions of reachable sets. The dual norm of the costate, $f^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f})$, is simply the support function of the sublevel set of the unit cost at time $t$, and $\max_{t\in[t_{0},t_{f}]}f^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f})$ is nothing but the support function of the pseudostate reachable set under unit cost [20].
@@
Let $\bm{\lambda}_{f}^{*}$ be the optimal solution to Problem 7 and $t^{*}=\{t\in[t_{0},t_{f}]\mid f^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}^{*})=1\}$. The set $t^{*}$ represents the optimal impulse times to the system. Neustadt [6] showed that the number of impulse times is at most $n_{x}$, and that selecting $n_{x}$ components of $t^{*}$ recovers the optimal semi-infinite variables. From these times, the optimal control can be constructed as $\bm{u}(t)=\sum_{t_{k}\in t^{*}}\bm{v}_{k}\delta(t-t_{k})$ where $\{\bm{v}_{k}\}_{k=1}^{n}$ satisfies
@@ III Problem Formulation
The novel contribution of this paper is the inclusion of input magnitude constraints on the control input. The consideration of input magnitude constraints is non-trivial, since simply enforcing that the norm of the control input be below a given value naturally breaks the impulsivity of the solution. Moreover, enforcing that any given impulse be below some impulse bound does not necessarily preserve the discretizability of the semi-infinite program.
@@
Find $\bm{u}(\cdot)$ that solves the primal optimization problem $\mathcal{P}=$
@@
$\int_{T_{k}}f_{k}(\bm{u}(t))\text{d}t\leq\Delta V_{max,k}\quad\forall k\in\mathbb{Z}_{1}^{N},$
where $f_{k}(\cdot)$ are norms, $\Delta V_{max,k}$ are fixed impulse magnitude limits, and $T_{k}$ represent ordered control input time windows that satisfy
$\bigcup_{k=1}^{N}T_{k}\subseteq[t_{0},t_{f}],\ T_{i}\cap T_{j}=\emptyset\ \forall i\neq j.$ (10)
@@ III-A Dual Formulation
To show that Problem 10 is amenable to being solved through semi-infinite programming, the dual form can be derived in a similar manner to Problem 1. The Lagrangian of Equation 9 is
$\begin{gathered}\mathcal{L}(\bm{u}(\cdot),\bm{\lambda}_{f},\{\sigma_{k}\}_{k=1}^{N})=\bm{\lambda}_{f}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})\\ +\sum_{k=1}^{N}\int_{T_{k}}(1+\sigma_{k})f_{k}(\bm{u}(t))-\bm{\lambda}_{f}^{\intercal}\Phi(t_{f},t)B(t)\bm{u}(t)\text{d}t\\ -\Delta V_{max,k}\sigma_{k},\end{gathered}$ (11)
and its dual function is given by
$\begin{gathered}g(\bm{\lambda}_{f},\{\sigma_{k}\}_{k=1}^{N})=\inf_{\bm{u}(\cdot)}\mathcal{L}(\bm{u}(\cdot),\bm{\lambda}_{f},\{\sigma_{k}\}_{k=1}^{N})\\ =\begin{cases}\bm{\lambda_{f}}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})-\sum_{k=1}^{N}\Delta V_{max,k}\sigma_{k},\\ \text{if}\ f_{k}^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t)f,t)\bm{\lambda}_{f})\leq 1\ \forall t\in T_{k},\forall k\in\mathbb{Z}_{1}^{N},\\ -\infty,\text{otherwise}\end{cases}\end{gathered}$ (12)
@@
Find $\bm{\lambda}_{f},\{\sigma_{k}\}_{k=1}^{N}$ to solve $\mathcal{D}=$
$\max_{\bm{\lambda}_{f},\bm{\sigma}\geq 0}$ $\bm{\lambda}_{f}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})-\sum_{k=1}^{N}\Delta V_{max,k}\sigma_{k}$ (13)
@@
This problem is indeed familiar to the unconstrained form of Problem 7. The problem consists still of maximizing an affine function subject to a semi-infinite constraint over the dual norm of the costate. Moreover, the problem has been reduced from an optimization over an infinite function space in the primal to an optimization over $n_{x}+N$ variables with $N$ semi-infinite constraints in the dual. This dual form permits the dual norm violation of $1$, but penalizes this violation in the cost weighted by the associated maximum impulse magnitude.
Several theorems are now presented to show that the dual form retains several desirable properties.
Theorem III.1
Problem 13 is strongly dual to the primal Problem 10, i.e. $\mathcal{D}=\mathcal{P}$.
@@
Evidently, Problem 13 satisfies Slater’s condition for semi-infinite programs (see Definition 2.1 from [23]) since there is a feasible interior point, i.e. $\exists\bm{\lambda}_{f},\bm{\sigma}>0,\text{s.t.}f_{k}^{\circ}(B^{\intercal}(t)\Phi(t_{f},t)\bm{\lambda}_{f})<1+\sigma_{k}\ \forall t\in T_{k},\forall k\in\mathbb{Z}_{1}^{N}$. From Theorem 2.3 of [23], this implies that strong duality holds for the semi-infinite program. ∎
This result is important for establishing feasibility of the primal problem, as it implies that when the dual problem is unbounded above ($\mathcal{D}=\infty$), then the primal problem is infeasible, and no input exists to steer the dynamical system between the boundary conditions subject to the input magnitude constraints.
Theorem III.2
The semi-infinite dual Problem 13 is reducible [23], i.e. there exists a finite subset of times $\{t\}\triangleq\{t_{1},\dots,t_{M}\}\subset\bigcup_{k=1}^{N}T_{k}$ such that
$\max_{\bm{\lambda}_{f},\bm{\sigma}\geq 0}$ $\bm{\lambda}_{f}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})-\sum_{k=1}^{N}\Delta V_{max,k}\sigma_{k}$ (14)
$\forall t\in T_{k}\cap\{t\},\ \forall k\in\mathbb{Z}_{1}^{N}$
provides an equivalent solution to Problem 13.
@@
This property is important in both the theoretical and algorithmic sense. With respect to the former, it shows that the optimal solution still takes the form of a sum of delta functions, meaning the solution structure is preserved under the formulation of Problem 10. With respect to the latter, discretization techniques can be applied to the semi-infinite dual problem to efficiently determine the solution of the primal problem.

### v2

@@ Abstract
This work presents a novel algorithm for impulsive optimal control of linear time-varying systems with the inclusion of input magnitude constraints. Impulsive optimal control problems, where the optimal input solution is a sum of delta functions, are typically formulated as an optimization over a normed function space subject to integral equality constraints and can be efficiently solved for linear time-varying systems in their dual formulation. In this dual setting, the problem takes the form of a semi-infinite program which is readily solvable in online scenarios for constructing maneuver plans. This work augments the approach with the inclusion of magnitude constraints on the input over time windows of interest, which is shown to preserve the impulsive nature of the optimal solution and enable efficient solution procedures via semi-infinite programming. The resulting algorithm is demonstrated on the highly relevant problem of relative motion control of spacecraft in Low Earth Orbit (LEO).
@@ I Introduction
The problem of minimizing the norm of a control input function space of a linear time-varying system with boundary conditions has received significant attention due to its practical application to many systems of interest. Such optimal control problems arise in many fields and are central to the study of spacecraft optimal control and trajectory optimization. Despite its widespread applicability, impulsive optimal control remains challenging for many applications of interest, which may involve complex state and input constraints and require efficient and reliable solution procedures. For example, in the context of spacecraft relative motion control, for which computational resources are limited, safety is essential, and algorithmic reliability is crucial, few existing approaches are adequate.
Approaches to solving impulsive optimal control problems are numerous, and include direct methods, such as sequential convex programming [18], indirect methods, leveraging moment problems [19], reachable set theory [10], or primer-vector theory [20], and other techniques, such as analytical methods [7]. In direct methods, the dynamics and cost are discretized over a time grid of interest and solved as a finite-dimensional optimization problem. Despite widespread utility, these techniques do not explicitly exploit solution sparsity and therefore tend to introduce many optimization variables or can produce suboptimal solutions in the case of coarse gridding.
Indirect methods, in comparison, involve formulating optimality conditions on the continuous-time problem, and include moment-based formulations [19], reachable set theory [10], and primer-vector theory approaches [16], which provide different perspectives to the problem and yet all arrive at equivalent optimality conditions.
@@
Early algorithmic frameworks for leveraging the optimality conditions derived from these indirect methods are proposed in Gilbert [10], which employed an early numerical algorithm based on quadratic programming [5] to efficiently solve impulsive rendezvous problems. Other approaches involve simplified techniques to solve the necessary conditions to suboptimality [6] or applying mixed iterative schemes to solving the polynomial necessary and sufficient conditions of optimality [3]. In 2013, Arzelier [4] connected the moment problem theorized by Neustadt [19] to a semi-infinite programming convex optimization scheme leveraging exchange methods [21] to produce an efficient, convergent, and reliable strategy to solving the linearized impulsive optimal control problem. This semi-infinite programming approach has served as the foundation for subsequent impulsive optimal control studies [14, 13] and has heritage on CubeSat missions such as VISORS [11].
@@
Despite the practicality of these theoretical and algorithmic frameworks, the applicability of indirect methods to constrained problems remains limited. With respect to spacecraft rendezvous, a major constraint of interest involves bounding the maximum impulse that a spacecraft can execute at a given time. All satellites, such as those of the VISORS mission, have propulsion systems that naturally constrain the amount of $\Delta V$ that can be applied at a given time, necessitating the use of maneuver splitting algorithms or other heuristic approaches [11] which lead to suboptimal maneuver plans or increased computational burden. More broadly, it is often practical to distribute $\Delta V$ expenditure for spacecraft relative motion irrespective of propulsion constraints to reduce the aggressiveness of maneuver plans.
@@
To address this theoretical and algorithmic gap, this paper introduces impulse magnitude constraints by applying inequality constraints to the integral of the input norm over specified time windows. It is shown that the resulting formulation preserves the impulsive quality of the solution and can be easily extended to semi-infinite programming techniques to yield an efficient and reliable solution procedure. The algorithm is demonstrated for several optimal control problems of interest in spacecraft relative motion control. In contrast to other formulations of input-constrained optimal control [1, 23], which are ubiquitous in aerospace literature, the formulation presented here is unique in that it preserves solution impulsivity, allowing for both improved algorithmic procedures and more desirable maneuver plans.
@@ II Problem Background
where $\bm{x}_{0}\in\mathbb{R}^{n_{x}}$ and $\bm{x}_{f}\in\mathbb{R}^{n_{x}}$ are fixed initial and terminal constraints. The set of input functions that satisfy the system dynamics and boundary conditions can be expressed as an equality constraint in terms of the state transition matrix via
@@
Problem 1 is a moment problem of the same form given in Neustadt [19], whose optimal solution is a sum of delta functions, i.e. the input is an atomic measure. In its primal form, this problem is difficult to solve due to its infinite dimensional nature. However, in the dual setting, the problem is conveniently reformulated as the following convex semi-infinite program
@@
$\sup_{\bm{\lambda}_{f}}$ $\bm{\lambda}_{f}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})$ (5)
@@
where the optimization variable $\bm{\lambda}_{f}$ is a Lagrange multiplier and $f^{\circ}(\cdot):\mathbb{R}^{n_{u}}\to\mathbb{R}^{+}$ is the dual norm of $f$, i.e. $f^{\circ}(\bm{y})=\sup_{f(\bm{x})\leq 1}\bm{x}^{\intercal}\bm{y}$. For many relevant norms, e.g., $l_{p}$ norms, the dual norm is analytically available ($\|\cdot\|_{p}^{\circ}=\|\cdot\|_{q},1/p+1/q=1$, see also [14]).
In contrast to the infinite dimensional primal problem, the dual problem is an optimization problem over a finite number of variables subject to an infinite constraint space. The dual form of this problem is in fact easier to solve in this form, as Arzelier [4] leveraged, through exchange methods for semi-infinite programs [21, 12].
The moment problem is connected to both primer-vector theory and reachable set theory approaches through its dual form. Indeed, $\bm{\lambda}_{v}(t)\triangleq B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}$ is in fact the classic primer-vector [16], with the Lagrange multiplier $\bm{\lambda}_{f}$ being the costate at the terminal time, and the semi-infinite constraint in Problem 5 is equivalent to that derived in Prussing [20]. From the perspective of reachable set theory [10], the dual norm of the costate, $f^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f})$, is simply the support function of the sublevel set of the unit cost at time $t$, and $\max_{t\in[t_{0},t_{f}]}f^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f})$ is nothing but the support function of the pseudostate reachable set under unit cost [14].
@@
Let $\bm{\lambda}_{f}^{*}$ be the optimal solution to Problem 5 and $t^{*}=\{t\in[t_{0},t_{f}]\mid f^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}^{*})=1\}$. The set $t^{*}$ represents the optimal impulse times to the system. Neustadt [19] showed that the number of impulse times is at most $n_{x}$, and that selecting $n_{x}$ components of $t^{*}$ recovers the optimal semi-infinite variables. From these times, the optimal control can be constructed as $\bm{u}(t)=\sum_{t_{k}\in t^{*}}\bm{v}_{k}\delta(t-t_{k})$ where $\{\bm{v}_{k}\}_{k=1}^{n}$ satisfies
@@ III Problem Formulation
The novel contribution of this paper is the inclusion of input magnitude constraints on the control input. The consideration of input magnitude constraints is nontrivial, since bounded control magnitudes, as are typically enforced in aerospace literature for continuous thrust formulations [1, 23], prevents the optimal solution from concentrating mass at discrete time points. As will be shown, the formulation presented here maintains impulsivity of the input profile, which provides both algorithmic benefits and a desirable input structure for spacecraft which are often modeled as impulsively actuated.
@@
Find $\bm{u}(\cdot)$ that solves the primal optimization problem $\mathcal{P}(\{T_{k}\}_{k=1}^{N})=$
@@
$\int_{T_{k}}f_{k}(\bm{u}(t))\text{d}t\leq\Delta V_{k}\quad\forall k\in\mathbb{Z}_{1}^{N},$
where $f_{k}(\cdot)$ are norms, $\Delta V_{k}$ are fixed impulse magnitude limits, and $T_{k}$ represent ordered control input time windows that satisfy $\bigcup_{k=1}^{N}T_{k}\subseteq[t_{0},t_{f}],\ T_{i}\cap T_{j}=\emptyset\ \forall i\neq j.$
@@ III-A Dual Formulation
To show that Problem 4 is amenable to being solved through semi-infinite programming, the dual form can be derived in a similar manner to that used for Problem 1. The Lagrangian of Equation 7 is
$\begin{gathered}\mathcal{L}(\bm{u}(\cdot),\bm{\lambda}_{f},\bm{\sigma})=\bm{\lambda}_{f}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})-\sum_{k=1}^{N}\Delta V_{k}\sigma_{k}\\ +\sum_{k=1}^{N}\int_{T_{k}}(1+\sigma_{k})f_{k}(\bm{u}(t))-\bm{\lambda}_{f}^{\intercal}\Phi(t_{f},t)B(t)\bm{u}(t)\text{d}t,\end{gathered}$ (8)
where $\bm{\lambda}_{f}$ and $\bm{\sigma}=(\sigma_{k})_{k=1}^{N}$ are Lagrange multipliers associated with the equality and inequality constraints of the primal in Problem 4 respectively. The associated dual function is given by
$\begin{gathered}g(\bm{\lambda}_{f},\{\sigma_{k}\}_{k=1}^{N})=\inf_{\bm{u}(\cdot)}\mathcal{L}(\bm{u}(\cdot),\bm{\lambda}_{f},\{\sigma_{k}\}_{k=1}^{N})\\ =\begin{cases}\bm{\lambda}_{f}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})-\sum_{k=1}^{N}\Delta V_{k}\sigma_{k},\\ \text{if}\ f_{k}^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f})\leq 1+\sigma_{k}\ \\ \forall t\in T_{k},\forall k\in\mathbb{Z}_{1}^{N};-\infty,\text{otherwise}\end{cases}\end{gathered}$ (9)
@@
Find $\bm{\lambda}_{f},\bm{\sigma}$ such that $\mathcal{D}(\{T_{k}\}_{k=1}^{N})=$
$\sup_{\bm{\lambda}_{f},\bm{\sigma}\geq 0}$ $\bm{\lambda}_{f}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})-\sum_{k=1}^{N}\Delta V_{k}\sigma_{k}$ (10)
@@
Similar to Problem 5, the problem has been reduced from an optimization over an infinite function space in the primal to a convex semi-infinite optimization problem over $n_{x}+N$ variables.
Lemma 1
Problem 10 is strongly dual to the primal Problem 4, i.e. $\mathcal{D}=\mathcal{P}$.
@@
Evidently, Problem 10 satisfies Slater’s condition for semi-infinite programs (see Definition 2.1 from [22]) since there is a feasible interior point, i.e. $\exists\bm{\lambda}_{f},\bm{\sigma}>0,\text{s.t.}f_{k}^{\circ}(B^{\intercal}(t)\Phi(t_{f},t)\bm{\lambda}_{f})<1+\sigma_{k}\ \forall t\in T_{k},\forall k\in\mathbb{Z}_{1}^{N}$. From Theorem 2.3 of [22], this implies that strong duality holds for the semi-infinite program. ∎
Lemma 2
The semi-infinite dual Problem 10 is reducible [22], i.e., there exists a finite subset of times $\{t\}\triangleq\{t_{1},\dots,t_{M}\}\subset\bigcup_{k=1}^{N}T_{k}$ such that $\mathcal{D}(\{t\})=\mathcal{D}(\{T_{k}\}_{k=1}^{N})$.
@@
Lemmas 1 and 2 extend several important properties to the proposed constrained problem. Namely, unboundedness of the dual problem implies infeasibility of the primal, the optimal solution retains impulsivity (i.e., $\bm{u}(t)=0$ a.e.), and exchange methods [21] are valid for solving Problem 10.